A conductor of length 0.4 m is moving with a speed of 7 m/s perpendicular to a magnetic field of intensity 0.9 Wb/m2. The induced e.m.f. across the conductor is


        (a)    1.26 V


        (b)    2.52 V


        (c)    5.04 V


        (d)    252 V


Grade:10

1 Answers

Simran Bhatia
10 years ago

(b)

Length of conductor () = 0.4 m;

Speed (v) = 7 m/s and magnetic field (B) = 0.9 Wb/m2.

Induced e.m.f. (V) = B/v sin B = 0.9 × 004 × 7 × sin 90° = 2.52 V.